A child has a piggy bank that contains only quarters and dimes. All together there are 27 coins in the bank and the total value of the coins is $4.50. How many quarters and how many dimes are in the piggy bank?
Answer by Alan3354(69443)

A child has a piggy bank that contains only quarters and dimes. All together there are 27 coins in the bank and the total value of the coins is $4.50. How many quarters and how many dimes are in the piggy bank?
----------
d + q = 27
10d + 25q = 450
--
Sub for d in the 2nd eqn and find q, or vice versa.
================
Or,
27 dimes = 270
Each dime replaced by a quarter adds 15 cents.
450 - 270 = 180
180/15 = 12 quarters
15 dimes.